OpenStudy (anonymous):

use the law of cosines to find the value of 2⋅3⋅4 cosθ

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@taramgrant0543664

OpenStudy (anonymous):

what do you mean by 2⋅3⋅4 cosθ You mean \[\cos 2\theta\quad,\cos3\theta,\quad \cos3\theta\]?

OpenStudy (taramgrant0543664):

Do you have a triangle that corresponds to this question?

OpenStudy (anonymous):